ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 172, 2009

Ben David, N. and Y. Ginosar On groups of central type, nondegenerate and bijective cohomology classes

317

Bourgain, J. On the distribution of the residues of small multiplicative subgroups of 𝔽_{p}

61

Costello, K. P. Balancing Gaussian vectors  145

Danilenko, A. I. and D. J. Rudolph Conditional entropy theory in infinite measure and a question of Krengel

93

Efrat, I. Valuations and diameters of Milnor K-rings  75

Ginosar, Y. See N. Ben David Ginzburg, D. Branching rules obtained from explicit correspondences

of automorphic representations

29

Girard, B. A new upper bound for the cross number of finite abelian groups

253

Han, B. and Z. Shen Characterization of Sobolev spaces of arbitrary smoothness using nonstationary tight wavelet frames

371

Hong, S. and S. Shin The entropies and periods of renewal systems  9

Horák, J. and M. Lucia A Minimax Theorem in the presence of unbounded Palais-Smale sequences

125

Isaev, A. V. and N. G. Kruzhilin Proper actions of Lie groups of dimension n^{2} + 1 on n-dimensional complex manifolds

193

Jaikin-Zapirain, A., M. F. Newman and E. A. O’Brien On p-groups having the minimal number of conjugacy classes of maximal size

119

Jaworski, W. On contraction groups of automorphisms of totally disconnected locally compact groups

1

Kim, J. H. On positive quaternionic Kähler manifolds with certain symmetry rank

157

Krieger, F. Minimal systems of arbitrary large mean topological dimension

425

Krushkal, S. Extremal problems for Fredholm eigenvalues  279

Kruzhilin, N. G. See A. V. Isaev Lucia, M. See J. Horák Mudrov, A. and V. Ostapenko Quantization of orbit bundles in

𝔤𝔩_{n}^{∗}(ℂ) 399

Newman, M. F. See A. Jaikin-Zapirain O’Brien, E. A. See A. Jaikin-Zapirain Ostapenko, V. See A. Mudrov Pinchasi, R. and G. Rote On the maximum size of an anti-chain of

linearly separable sets and convex pseudo-discs

337

Robertson, D. and A. Sims Simplicity of C^{∗}-algebras associated to row-finite locally convex higher-rank graphs

171

Rote, G. See R. Pinchasi Rudolph, D. J. See A. I. Danilenko Shen, Z. See B. Han Shin, S. See S. Hong Sims, A. See D. Robertson

Trahtman, A. N. The road coloring problem  51
Wilkens, B. p-groups without noncharacteristic normal subgroups  357

Xu, M. and X. Yang Remark on solvability of P -Laplacian equations in large dimension

349

Yang, S. Elliptic elements in Möbius groups  309

Yang, X. See M. Xu